OpenStudy (anonymous):

Write the equaiton of the line that passes through (3, 5) and (6, 6) in slope intercept form.

OpenStudy (wach):

Alrighty, so slope intercept form is basically : y=mx+b, right? So we definitely need to find the slope, which is basically the 'average' between points: (y2-y1) / (x2-x1) This way, we're finding the average change between the points. So: 6-5 / 6-3= 1 / 3 So our slope is 1/3 We could find the y-intercept, but there's an easier way: we can put this in point-slope form where y - y1 = m (x - x1) let's plug in (3,5) y - 5 = 1/3 (x - 3) y - 5 = 1/3x - 1 y = 1/3x +4

OpenStudy (anonymous):

You are the best!!!!!!!!

OpenStudy (wach):

No problem!